# The unified cross-disciplinary model of the operation of neurons

### v2

@@ Abstract
Physics perfectly describes neuronal operation, provided that we take into account that biology uses slow positive ions instead of electrons as charge carriers and remove untested ad-hoc hypotheses that contradict science’s first principles. We also incorporate recent experimental discoveries into the outdated classic theoretical description. Lipid mechanisms are really very important for cellular biology, but they are certainly not suitable for describing the phenomena we discuss. We introduce the correct physical model, significantly enhancing the classic Hodgkin and Huxley (HH) model; furthermore, the fundamentally bio-electrically triggered operation causes changes in electrical, mechanical, and thermodynamic properties of living matter. We derive the resting potential from first principles of science, showing that it has no relation to an ad-hoc linear combination of mobilities or reversal potentials as the Goldman-Hodgkin-Katz (GHK) equation claimed. Furthermore, we derive an "equivalent thermodynamic electric field" that enables discussion of, among others, the operation of ion channels, their ion selectivity, and voltage sensing. We demonstrate that a simple electrical-thermodynamic control circuit regulates neuronal operation, setting and maintaining a stable resting potential and handling an unstable transient process known as the Action Potential (AP). Its setpoint entirely defines the resting potential, explaining its robustness during growth and evolution. Our cross-disciplinary approach naturally fuses the electrical and mechanical/thermodynamic description of neuronal operation, resolves the decades-old mystery of "heat absorption" and "leakage current" (with their far-reaching consequences), and derives the thermodynamic description of neural computing. We defy that science cannot describe life.
@@ 1 Introduction
The excellent textbook [1] separates the neuron membrane’s states into resting and transient states. The statement that "the resting membrane potential results from the separation of charge across the cell membrane" is only half the truth; here we tell the second half. We refine their statement "by discussing how resting ion channels establish and maintain the resting potential"  [1], page 126. In the resting state, two effects are considered. On the one side, the potential is created by a static charge-up: "the cell membrane potential results from the separation of net positive and net negative charges on either side of the membrane". However, as we discuss, it is only a small part of the potential: the polarization of the electrolyte’s dipoles contributes the rest. (For the sake of clarity, both mechanisms are known in physics. The word ’polarization’ is reserved for rearranging charges in multiply-charged systems without the application of an external current. The changes in the membrane’s potential difference are a consequence of complete charge separation (and the movement of slow charges), even if physiology refers to it as polarization. External currents contribute to the changes of potential.) On the other hand, it was claimed that "the resting membrane potential is the result of the [low intensity] passive flux of individual ion species through several classes of resting channels" [1], although it was question-marked whether the membrane permeability is a primary property for the membrane potential generation, for a review, see [2]. The claim is rooted in the hypothesis and the measurements by Hodgkin and Huxley [3], as we discuss in section 2.6.4. Their numeric figure was correct, but they mistakenly interpreted the voltage they measured as being produced by a "leakage current" on the membrane’s parallel resistance. Furthermore, their interpretation leads to causality issues [4]. Paper [5] stated correctly "that separation of charges produces a potential" and derived the thermodynamic part in two different ways, but did not account for the electric potential. We derive how that electric potential is produced by separating charges and how its value can be calculated from first principles.
@@
Our results underpin that "the crucial system in biology isn’t a molecule or a molecular class whatsoever, but the interface created by biomolecules in water" [6]. More precisely, inseparable thermodynamic and electric processes keep the potential set by the inseparable electric and thermodynamic backbone defined by the lipid structure of the cell. It is one of the frequent cases when one effect implements a functionality (the backbone closely defines the frames of the operation), and the other (in this case, combined thermodynamics and electricity) corrects it when it implements a complex operational (dynamical) functionality: "stimulated phase transitions enable the phase-dependent processes to replace each other … one process to build and the other to correct" [7].
The book [1] asks the central questions, "How do ionic [i.e., electrical and chemical] gradients contribute to the resting membrane potential? What prevents the ionic gradients from dissipating by diffusion of ions across the membrane through the resting channels?" However, it leaves them essentially open by giving only a qualitative answer, as it discusses only membrane permeability without explaining how the resting potential is created and regulated through charging and polarization. We demonstrate in section 3 that classical electric methods enable us to calculate the membrane’s potential resulting from charge separation and polarization. However, one can make some physically plausible assumptions to provide numerical figures. The electric processes are inseparable from the corresponding thermodynamical processes [8], which we discuss by deriving an "equivalent thermodynamical electric field" in section 2.3. The other minor effect is a low-intensity "passive flux" (material transport) produced by the "resting channels". They are ion channels, not necessarily only across the membrane; they also include the ion channels in the Axon Initial Segment (AIS) [9, 10, 11, 12, 13]. The "resting current" is about two orders of magnitude lower [14, 15] than it was assumed at the time of writing the book; correspondingly, its role is much smaller in the transient state, and it even drastically changes the electric model used in producing an AP which we discuss elsewhere [16].
@@ 2 Membrane’s electricity
The importance of the subject was correctly estimated decades ago [17]: "For the past forty years, our understanding and our methods of studying the biophysics of excitable membranes have been significantly influenced by the landmark work of HH. Their work has had a far-reaching impact on many different life science sub-disciplines where concepts of cell biology have come to be important. …The ionic currents and electrical signals generated by neuronal membranes are of obvious importance in the nervous system." To fix the flawed and incomplete hypotheses has a similar importance.
It was already stated that the membrane’s resting potential is, of course, of electrostatic origin [18, 19, 20, 2, 21]. The closed volume of the neurons drastically differs from the infinite space: the charges repel each other, but since they experience the counterforce of the limiting surface (the membrane), a static mechanical pressure and a shock wave are also created, that naturally connects the electric and mechanical effects, as well as the electrical, mechanical, and thermodynamic theories of operation. We must discuss them in a non-disciplinary way because the participating ions belong simultaneously to the disciplines of electricity and thermodynamics; furthermore, we must cross the borderline between particle-based and continuous points of view of nature [8]. Furthermore, the features of living matter may change while being measured in physics laboratories as usual: "the construction is different from anything we have yet tested in the physical laboratory" [22].
In the absence of external influences (which includes the lack of a separating membrane), both gradients are balanced and are at zero. When this state is perturbed, the system attempts to find a new balanced state, using temporal processes. In this scenario, the ’carrier’ - the ion - can be influenced by two interactions. We discussed elsewhere that those phenomena in living matter need deriving and applying ’non-ordinary’ laws of science [8] , and the time course of those processes can be described by the laws of motion of biology. Fundamentally, due to the mixing of interaction speeds, considering only one of the interactions (the disciplinary way) leads to incorrect conclusions and theoretical interpretations.
@@ 2.1 Condenser
Biology represents a complex case where the phenomena cannot be reasonably reduced to a single discipline, as classic inanimate science does when deriving its laws for "material points", as A. Einstein coined, abstracted to either mass or charge, but not both. Cell biochemistry sufficiently well describes the lipid bilayers that constitute the membrane [23, 24], but does not explain that the negative and positive ends of lipids in an electrolytic solution attract and absorb ions, forming a charged layer on both surfaces of the membrane. In this way, the membrane in electrolytes acts as a condenser. However, this condenser comprises two charged sheets on the surfaces of the membrane with ill-defined parameters. Their boundaries, thicknesses, and compositions depend on the operating state (as opposed to conducting plates with well-defined sizes and boundaries). This condenser is not ideal: the charged sheet comprises ions with attached aqueous constituents from the electrolytes, so its thickness is far from the idealized zero-thickness known from classical electricity. Furthermore, in its wall, it inseparably comprises ion channels (from an electric point of view: resistors), and the condenser plates are in polarizable electrolyte solutions.
Cell science does not establish its electrical phenomena; instead, it attempts to elucidate complex lipid activities that explain charge transfer, resting and transient potential, and the entire biological operation using mathematical formulas without a realistic physical background. Electrophysiology combines currents of electrons and ions, and uses Ohm’s Law for its admittedly non-Ohmic systems. It introduces foreign (clamping) currents into biological systems, attributing their effect to a changed conductance of the membrane without explaining how the voltage and current can be independent of charge carriers.

### v3

# The unified non-disciplinary model of the operation of neurons

@@ Abstract
The paper introduces the unified, non-disciplinary theory of neuronal operation that explains all known experimental observations without contradictions. The competing electrical and thermodynamic disciplinary theories mutually cannot explain some observations of neuronal operation, simply because some quantities fall outside of the scope of the given discipline. Although the first principles of science are identical in the inanimate and life sciences, the abstractions and approximations that underlie concrete laws differ, so the actual laws depend on the environment. We challenge the universality of protein mechanisms for describing neuronal operation and show that, by using the appropriate approximations, the non-disciplinary cooperation of thermodynamics and electricity perfectly describes processes in living matter. We show that biology works on the boundary of the micro and macro views of physics, and in some cases, those views must be bridged. Most importantly, we explicitly consider the discrete nature of biological current, that the charge carriers are slow and they repel each other. That feature, overlooked in both mentioned theories, enables us to interpret the electrical and thermodynamic behavior of important biological objects and processes. The speed of biological currents changes by several orders of magnitude. It keeps the electrolyte in continuous movement, making some small regions of the volume responsible for the observed action potential. The cross-disciplinary model enables us to derive the correct electrical equations and thermodynamic description of the action potential. In a way demonstrating the need for the non-disciplinary discussion, we derive the thermodynamic Carnot-cycle of AP from measuring electrical parameters, calculate its energy consumption and efficiency, interpret the neuron’s partially reversible operation and the long-sought solution to the heat emission/absorption issue. We defy the claim that science cannot describe life: the underlying physics constrains biological phenomena, including cognitive ones.
@@ 1 Introduction
The dynamic operation of individual neurons, their connections, and higher-level organizations, connections, the brain with its information processing capability, and finally, the mind with its conscience and behavior, are still among the big mysteries of science: at which point the non-living matter becomes a living one and at which point the living matter becomes intelligent and conscious. Moreover, whether and how science can handle all this stuff. We desperately need a new understanding. In neuron models, dynamics refers to how the neuron’s state variables (primarily membrane potential and ion channel properties) change over time in response to internal processes and external inputs. However, unlike in our approach, the usual processes use a time parameter in empirical or mathematical functions; see, for example,  [1]. We use a genuine dynamics, based on physics, in the sense that Newton introduced his laws of motion to describe response of neurons.
The "great journey into the unknown [understanding the brain]" [2] must begin at a much lower level: revisiting the fundamental phenomena, disciplines, laws, interactions, abstractions, omissions, and testing methods of science. Research must build on classical science, but be reinterpreted for living matter. There is no independent ’life science’, there is only science. It is based on the same ’first principles’ but using different abstractions and approximations for living and non-living matters, and having the appropriate relations between them. "There is a clear need for a tighter and more carefully managed integration and realignment of the work" [3]. Without aligning the knowledge elements along the first principles, the "integration between data and knowledge from different disciplines", lacks integration. However, even the review of the targeted Human Brain Project (HBP) summarized that "HBP is not developing with the expected level of integration and the project controls in place are not adequate to achieve this aim." [3]. We arrived at the boundaries of classical science disciplines and are moving now through terra incognita.
The hype in the newly launched projects is excessive. For example, the newly (at the end of 2025) launched “https://brainminds.jp/en Brain/MINDS 2.0” program in Japan was launched using only mathematical models, without targeting the understanding of the underlying physical processes. "A biological model is often understood to be simply a diagram depicting the interrelationships of various (sub)systems in a process, whereas a physical model is expected to be a theoretical description of a process involving a number of equations of motion stemming from the first principles (if possible), testable against a range of tunable experimental conditions. It must lead to a quantitative prediction and not simply reproduce already known results" [4].
Our model is undoubtedly a physical model, in the above sense. We provide a holistic, multi-level view of neuronal operation and describe (at least semi-)quantitatively the actual physical processes; this way, we achieve the promised ’new understanding’ (and the need for rewriting books in neuroscience). Although the classical neuron model aims to be an electrical model, it violates fundamental laws of electricity and conservation laws; furthermore, it cannot interpret phenomena from other disciplines (thermodynamics, mechanics, optics). The thermodynamic model(s), on the other hand, cannot explain the electrical features of the observed processes. We prove that no compromise is needed: the pressure and the voltage measurements observe the same physical process by different disciplinary means. Neuroscience forgot that neural currents consist of ions, which repel each other in a confined volume that contains an object with static charge; furthermore, that ions are charged and have mass, both of which are abstractions from different classical disciplines.
Research already provided the detailed description of a correct ‘net electrical’ neuron model [5], the laws of motion for describing biology [6], and now provides a correct non-disciplinary model (promising a unified and scientifically complete electrical/thermodynamic model). The deeper reason for the apparent conflict between the two competing theoretical descriptions is science’s disciplinarity, as we discuss it separately in Appendix A, since it is somewhat philosophical and touches not only the fundamental concepts of physics but also the way of scientific thinking. Shortly, the disciplinary separation proved to be successful for most phenomena of inanimate science. Electrolytes are an exception. Thermodynamics (due to excluding a long-range interaction between its particles), excludes electric charge (and related quantities) from the quantities it studies. Electricity (due to implying ’electron cloud’ as a charge transmission method) similarly excluded that long-range interaction may exist between its discrete charge carriers; that is, that they may have thermodynamic features based on the Vlasov equations. This way, electricity excludes mass (and related quantities, such as heat) from the quantities it studies. Neither of the two disciplines, alone, can completely describe electrolytes (and so: biological systems): some of the system’s quantities remain outside the scope of the given discipline. The measurements are not complete: some quantities are not measured, so ’miracles’ can occur. The force analysis does not show a driving force, so some "new force or whatnot" [7] must be assumed (in physiology, "protein mechanisms"). Our non-disciplinary approach, however, can successfully describe the operation of a neuron, without a miracle.
The structure of the paper is as follows. Section 2 presents the state-of-the-art of theoretical neuron modeling. A strongly simplified discussion in section 3, without mathematical, physical, and electrical details, presents the unified model. Electricity of ions and its application to the case of neurons is the subject of section 4, which includes the specifics of ion currents, the electrical features and operation of biological objects, and, first of all, the thermodynamics applied to neurons. Section 5 presents the physical mechanisms that, instead of a magic protein mechanism, define and maintain the membrane’s potential in resting and transient states. The importance of the action potential deserved a separate section 6, where the physical processes, their correct "net electrical" model, with the related mathematics, are described. The discussion culminates in section 7, which on one side derives the thermodynamic parameters describing producing Action Potential (AP) and their long sought calculation method, on the other provides evidence that only the cooperation of the competing disciplines can succeed: the thermodynamic parameters can be calculated by measuring parameters of electrical processes. The numerical estimations, serving as the base for some statements of the paper, are detailed in Appendix B.
2 State-of-the-art
Today, two major branches of disciplinary theories compete for being applied to describe the neuronal operation [8, 9, 10, 11, 12, 13]. Those cited references also compare those theories. The electrical view essentially stems from the (Nobel Prize-winning) work of Hodgkin and Huxley [14]. They could not make a perfect job, mainly due to the lack of discoveries made several decades later, including ours, about handling the finite speed of the biological ionic currents (and, due to that, the dynamic features), which drastically change their conclusions. In contrast with their empirical description, which delivered mathematically formulated measured observations, our discussion, although it follows essentially the same principles, reinterprets and pinpoints the used fundamental terms, sets up a physical model, and explains the physically underpinned processes. “However, the theory could not explain the physical phenomena such as reversible heat changes, density changes, and geometrical changes observed in the experiments” [11]. The thermodynamic view roots in the (possibly Nobel-prize-winning) idea of Heimburg and Jackson [15]. They proposed that the action potential is essentially a sound wave (a soliton). “However, there are several other questions that this has to answer like ion flow involvement in nerve signal propagation as stated by the Hodgkin and Huxley (HH) model and also the faster propagation in myelinated nerves than in unmyelinated” [11]. If we consider that the ion flow means charge propagation in a membrane tube where the specific capacity depends on the thickness of the wall (the myelin sheath) [16], the faster propagation is not mysterious anymore. Given that our (non-disciplinary) model natively connects charge and mass of ions [6], and it can explain the missing phenomena.
The excellent textbook [17] separates the neuron membrane’s states into resting and transient states. The statement that "the resting membrane potential results from the separation of charge across the cell membrane" is only half the truth; we tell the second half in section 4. We refine their statement "by discussing how resting ion channels establish and maintain the resting potential" [17], page 126. As we show in section 5, two different physical mechanisms operate in the resting and the transient states. A primary reason for misguiding neurophysiology was projecting the mechanism of the resting state to the transient state.
@@ 2 State-of-the-art
The book [17] asks the central questions, "How do ionic [i.e., electrical and chemical] gradients contribute to the resting membrane potential? What prevents the ionic gradients from dissipating by diffusion of ions across the membrane through the resting channels?" However, it leaves them essentially open by giving only a qualitative answer, discussing only membrane permeability without explaining how the resting potential is created and regulated through charging and polarization. We demonstrate in section 5 that classical methods of electricity enable us to calculate the membrane’s potential resulting from charge separation and polarization. Our results underpin that "the crucial system in biology isn’t a molecule or a molecular class whatsoever, but the interface created by biomolecules in water" [18]. However, we add that mainly physical processes and speed gradients control the processes. More precisely, inseparable thermodynamic and electrical processes keep the potential set by the inseparable electrical and thermodynamic backbone defined by the lipid/protein structure of the cell. It is one of the frequent cases when one effect implements a functionality (the backbone closely defines the frames of the operation), and the other (in this case, combined thermodynamics and electricity) corrects it when it implements a complex operational (dynamical) functionality: "stimulated phase transitions enable the phase-dependent processes to replace each other … one process to build and the other to correct" [19].
3 Operation in a nutshell
We have E. Schrödinger’s comments in mind when constructing the physical picture of neurons, that we must be cautious. We must also include the discoveries from the past seven decades, among others, that in the transient state, a neuron shall be modeled as a serial (i.e., differentiator-type) electrical oscillator, which has gated current inputs (instead of a parallel oscillator without gating inputs, which is a good model only in the resting state). We explicitly add that the charge carriers, the neuron works with, are ions (instead of electrons), and that the signal propagation mechanism is bioelectric (instead purely an electric field or mechanical wave propagation or "protein mechanisms"), in many cases even without having free carriers in the respective volume. (We do not consider the ideas that the associated mechanical, optical, etc. changes cause electrical phenomena. They reverse the true causality: they do not have a triggering cause; the view inherited from the narrative biology.) Electrochemical operation, especially that of the semipermeable membranes, requires special attention: charge carriers are "created" inside the biological matter. Omitting the precise physical background results in performing wrong measurements and the introduction of wrong concepts into biology.
In the non-disciplinary physical model (the ’abstract physical neuron’, as we call it), a neuron (highly simplified) is a sphere with two lipid layers (membrane), a similarly insulating output tube (axon), and an incompressible, partially ionized electrolyte fluid with different composition and concentration inside and outside. They contain ions, simple chemical molecules, and complex biological formations. These components may have gradients and move at continuously changing speeds that can vary by several orders of magnitude during operation. The lipid layer is partially permeable (controlled or uncontrolled) for the components above (channels, mainly ion channels).
The membrane represents a confined space in which ions behave differently from in an infinite space; furthermore, it significantly alters the properties of the few-nanometer-thick layer near the membrane. The elastic wall of the membrane is covered with ion layers, and therefore, a potential difference is created there (the resting potential). In this layer, the concentration and potential change significantly with distance from the membrane due to interactions between the membrane’s ion layers and the electrolyte. Ions sent by other neurons can enter the internal volume through controlled inputs (synapse). Between the membrane and the axon, there is a large amount of uncontrolled ion channels (Axon Initial Segment (AIS)). The conductance of the channels in the membrane’s wall is about 50 times smaller [20, 21] than that in the AIS.
The membrane functions as a Proportional-Integral-Derivative (PID) control circuit. Under a small external influence, a neuron remains in its resting state and regulates itself through non-controlled ion channels in its membrane. If the external influence exceeds a threshold, the controlled channels suddenly release a large amount of ions into the internal space, putting the neuron into a transient state. The membrane potential changes suddenly (within a few dozen nanoseconds), and, due to the mutual repulsion of ions, the pressure acting on the membrane wall and the ion concentration in the layer near the membrane also increase significantly.
Since the same physical action generates the voltage gradient and the concentration gradient (and, consequently, the pressure), they are inseparable and proportional to each other. Generating a voltage change (clamping, direct current, magnetic pulse) or a pressure change (ultrasound, mechanical, shock waves) on the membrane causes the other to change. Inputting ions (synaptic input, followed by rush-in) generates both changes. The control circuit tries to restore the resting state, during which the mentioned changes decrease proportionally together. The transient state of the neuron can, in principle, be described by any of the mentioned entities, but, as a practical matter, the effects of changes in the other entities must also be taken into account: no single discipline can describe the changes. The elastic membrane implements an almost critically damped vibration, but the "vibrating mass" and the "spring force" are constantly changing. In an electrical resonant circuit, the capacitor’s voltage and capacitance change due to the finite amount of charge stored on it.
The pressure wave caused by a force shock and the voltage wave caused by a voltage shock describe the same effect. In both cases, the other effect must be taken into account to some extent: an infinitesimal movement of the ion also involves changes in both pressure and voltage. Changes in pressure and charge alter the membrane’s thickness (and therefore its capacity), and both voltage and pressure can cause a structural change (a phase change, "melting") in the lipid chains that make up the membrane. In the case of the pressure wave, the speed is naturally finite (mass must be moved). For the electrical wave, there are laws regarding only instantaneous interaction; so, they must be adapted to slow ion currents based on physical assumptions. It is simpler to describe the generation of the action potential as an electrical oscillation (although slow currents’ propagation must be accounted for), and its propagation as a pressure wave (although this only represents well the wave front: the AP lasts as long as there is non-equilibrium charge in the neuron: the repulsion between the ions pushes the ions out of the closed space, so the intensity of the wave continuously decays). Hyperpolarization is described in the electrical view as a capacitive current in the sequential oscillator circuit, in the thermodynamic view as a suction effect occurring in the opposite phase of the membrane. In both cases, however, it must be taken into account that charge and mass are inseparable in ions, and that their cross-disciplinary effects significantly reduce the applicability of the disciplinary laws.
In simple words, a neuron combines different disciplines; it works like a two-tact biological internal combustion engine, and can be described theoretically as a special Carnot-type thermodynamic engine. The synaptic input currents operate the ignition. The influx of $Na^{+}$ ions acts like an explosion, producing a large pressure impulse due to electrical repulsion between ions; see Appendix B. The neuron’s volume remains practically unchanged (represents an iso-volume process). The longitudinal pressure wave vibrates the particles (including the dissociated ions). It is similar to a sound wave, with negligible material transport. The created offset potential moves the ions out of the neuron through the only available output channel: the axon.
In the first stroke, the pressure impulse invests energy into compressing the elastic membrane and starting a compression wave (a soliton). In the second stroke, a combined thermoelectrical/mechanical force (rather than some magic battery) pumps out ions through the ion channels in the AIS, which "resists" it, thereby generating a potential wave known as AP; well described by electricity. The force analysis shows that the pressure wave causes a longitudinal vibration, while the changed potential adds a potential-dependent longitudinal force component for moving the ions.
The compression produces heat, as evidenced by a rise in temperature. The expansion consumes heat, as evidenced by a decrease in temperature. These are well-understood processes in thermodynamics. However, they are still unexplained in biophysics, although they were observed seven decades ago [22]. The elastic membrane functions as a damped oscillator, which, in its negative amplitude phase, reverses the direction of electrolyte flow (and, consequently, the direction of the current carried by the ions). Correspondingly, one can observe, in addition to the electrical potential wave, mechanical, optical, and other features changing. In the discipline of electricity (when using the serial $RC$ model), it can be interpreted as a capacitive current, which physiology observes as hyperpolarization that inverts the polarity of the output voltage. (By mistake, physiology interprets the capacitive current that the outward current in the channel consists of $Na^{+}$ ions, and when the resultant potential changes its direction, the ion changes to $K^{+}$; a kind of alchemy.) In different disciplines, pressure and voltage represent the same phenomenon, using concepts from various disciplines. However, the speeds of those waves are strongly different: The AP as an electrical wave is a million times slower than the one described by the Maxwell laws, given that it is a side effect of the mechanical wave.
@@ 4 Membrane’s electricity
The importance of the subject was correctly estimated decades ago [23]: "For the past forty years, our understanding and our methods of studying the biophysics of excitable membranes have been significantly influenced by the landmark work of HH. Their work has had a far-reaching impact on many different life science sub-disciplines where concepts of cell biology have come to be important. The ionic currents and electrical signals generated by neuronal membranes are of obvious importance in the nervous system." To fix the flawed and incomplete hypotheses has a similar importance.
Concerning the origin of the membrane’s resting potential, we disagree that "the resting membrane potential is the result of the [low intensity] passive flux of individual ion species through several classes of resting channels" [17]. The claim was already question-marked whether the membrane permeability is a primary property for the membrane potential generation, for a review, see [24]. This fallacy is rooted in the hypothesis and the measurements by Hodgkin and Huxley [25], as discussed in section 4.5.4. Their numeric figure was correct, but they mistakenly interpreted (due to erroneously projecting the mechanism in the resting state to the transient state) the voltage they measured as being produced by a "leakage current" on the membrane’s parallel resistance. Furthermore, their interpretation leads to causality issues [12].
It was already stated that the membrane’s resting potential is, of course, of electrostatic origin [26, 27, 28, 24, 29]. Paper [30] stated correctly "that separation of charges produces a potential" and derived the thermodynamic part in two different ways, but did not account for the electrical potential. However, as discussed below, it is only a small part of the potential; the rest is contributed by the polarization of the electrolyte’s dipoles.
Biology represents a complex case where the phenomena cannot be reasonably reduced to a single discipline, as classic inanimate science does when deriving its laws for "material points", as A. Einstein coined, abstracted to either mass or charge, but not both. Moreover, the closed volume of the neurons drastically differs from the infinite space. The charges attract and repel each other, but since they experience the counterforce of the limiting surface (the membrane), a static mechanical pressure and a shock wave are also created, which naturally connect the electrical and mechanical effects, as well as the electrical, mechanical, and thermodynamic theories of operation. We must discuss them in a non-disciplinary way because the participating ions belong simultaneously to the disciplines of electricity and thermodynamics; furthermore, we must cross the borderline between particle-based and continuous points of view of nature [31]. Moreover, the features of living matter may change while being measured in physics laboratories as usual: "the construction is different from anything we have yet tested in the physical laboratory" [7].
When the balanced state is perturbed, the system attempts to find a new balanced state, using temporal processes. In this scenario, the ’carrier’ - the ion - can be influenced by two interactions. We discussed elsewhere that those phenomena in living matter need deriving and applying ’non-ordinary’ laws of science [31], and the time course of those processes can be described by the laws of motion of biology. Fundamentally, due to the mixing of interaction speeds, considering only one of the interactions (the disciplinary way) leads to incorrect conclusions and theoretical interpretations.
@@ 4.1 Condenser
Cell biochemistry sufficiently well describes the lipid bilayers that constitute the membrane [32, 33], but does not explain that the negative and positive ends of lipids constituting the membrane, in an electrolytic solution, attract and absorb ions, forming a charged layer on the surfaces of the membrane. In this way, the membrane in electrolytes acts as a condenser. However, this condenser comprises two charged sheets on the membrane surfaces, with ill-defined parameters. Their boundaries, thicknesses, and compositions depend on the operating state (unlike conducting plates with well-defined sizes and boundaries). This condenser is not ideal: the charged sheet comprises ions with attached aqueous constituents from the electrolyte, so its thickness is far from the ideal zero thickness of classical electricity. Furthermore, in its wall, it inseparably comprises ion channels (from an electrical point of view: resistors), and the condenser plates are in polarizable electrolyte solutions.
Cell science does not establish the membrane’s electrical phenomena; instead, it attempts to elucidate complex protein activities that explain charge transfer, resting and transient potentials, and explains the entire biological operation using mathematical formulas without a realistic physical background. Electrophysiology combines currents of electrons and ions, and uses Ohm’s Law for its admittedly non-Ohmic systems. It introduces foreign (clamping) currents into biological systems, attributing their effects to changes in membrane conductance without explaining how voltage and current can be independent of charge carriers.